# Causal Graph Neural Networks for Healthcare

### v2

@@ Abstract
Healthcare artificial intelligence systems routinely fail when deployed across institutions, with documented performance drops and perpetuation of discriminatory patterns embedded in historical data. This brittleness stems, in part, from learning statistical associations rather than causal mechanisms. Causal graph neural networks address this triple crisis of distribution shift, discrimination, and inscrutability by combining graph-based representations of biomedical data with causal inference principles to learn invariant mechanisms rather than spurious correlations. This Review examines methodological foundations spanning structural causal models, disentangled causal representation learning, and techniques for interventional prediction and counterfactual reasoning on graphs. We analyse applications demonstrating clinical value across psychiatric diagnosis through brain network analysis, cancer subtyping via multi-omics causal integration, continuous physiological monitoring with mechanistic interpretation, and drug recommendation correcting prescription bias. These advances establish foundations for patient-specific Causal Digital Twins, enabling in silico clinical experimentation, with integration of large language models for hypothesis generation and causal graph neural networks for mechanistic validation. Substantial barriers remain, including computational requirements precluding real-time deployment, validation challenges demanding multi-modal evidence triangulation beyond cross-validation, and risks of causal-washing where methods employ causal terminology without rigorous evidentiary support. We propose tiered frameworks distinguishing causally-inspired architectures from causally-validated discoveries and identify critical research priorities, including hierarchical causal discovery exploiting biological modularity, integration of diverse data modalities, and regulatory pathways for systems making causal rather than purely associational claims. Success requires balancing theoretical ambition with empirical humility, computational sophistication with clinical interpretability, and transformative vision with uncompromising validation standards.
@@ 1 Introduction
The convergence of causal inference with graph neural networks addresses this challenge. Yet, standard GNNs inherit and amplify the correlation-causation conflation, learning to exploit whatever statistical patterns maximise performance regardless of biological plausibility or causal validity [26, 27]. Recent theoretical advances in causal representation learning provide a path forward. By explicitly modelling causal structures within graph architectures, we can develop models that identify genuine therapeutic targets, predict treatment responses under interventions never observed in training data, and discover disease mechanisms rather than spurious correlations [28, 29]. This convergence is particularly timely given the four concurrent developments. First, the accumulation of multi-modal biomedical data, from single-cell sequencing that captures 20,000+ gene expression profiles to continuous physiological monitoring that generates terabytes of temporal dynamics, provides unprecedented resolution for causal discovery. Datasets now capture interventional outcomes and temporal dependencies that were previously unavailable [30, 31, 32]. Second, regulatory frameworks increasingly demand explainable AI in healthcare, which requires not merely predictions, but a mechanistic understanding that clinicians can verify against domain expertise and biological knowledge [33, 34]. Third, documented failures of correlation-based models under distribution shift have created urgency for methods achieving robust generalisation through causal invariance. Fourth, the integration of causal inference with fairness constraints enables rigorous identification and mitigation of discrimination transmitted along specific causal pathways, providing tools for distinguishing legitimate predictive factors from discriminatory proxies, a distinction central to anti-discrimination law yet absent from traditional machine learning approaches [13, 35].
@@
This review examines causality-inspired graph neural networks (CIGNNs) as a methodological framework for transforming AI in healthcare from reactive pattern recognition to proactive causal medicine. Unlike previous reviews that addressed GNNs in healthcare [36, 37] or causal inference in medicine [16, 38] separately, we focus specifically on their integration, methods in which both graph structure and causal reasoning prove essential for clinical impact. We analyse how causal principles enhance GNN architectures for healthcare tasks, moving beyond correlation to mechanism, and examine current applications that demonstrate clinical utility across psychiatric diagnosis, cancer treatment, physiological monitoring, and drug discovery. We also identify fundamental challenges in scaling these methods to real-world deployment, including computational complexity, validation without ground truth, fairness under distribution shift, and regulatory considerations. Our analysis reveals both transformative promise and substantial barriers: while CIGNNs achieve superior generalisation across healthcare institutions, identify biologically plausible mechanisms, and maintain fairness properties under distribution shift, they face computational requirements precluding real-time application, difficulty validating causal claims from observational data, and gaps between methodological sophistication and clinical interpretability.
@@
The result of this research is the development of patient-specific Causal Digital Twins, dynamic computational models built on causal GNN frameworks that enable clinicians to perform in silico experiments before clinical intervention. Imagine a clinician treating advanced cancer who could load the patient’s multi-omics profile, brain imaging, and clinical history into such a system, then simulate five drug combinations to predict effects on specific tumour pathways, toxicity risks, and progression-free survival, identifying optimal personalised therapy before administering a single dose. This vision, while ambitious, rests on methodological foundations now being established through causality-inspired architectures for interventional prediction, counterfactual generation, and robust multi-modal integration. Realising this potential requires not only technical innovation but fundamental rethinking of how we evaluate and deploy AI in medicine. We need to shift from predictive accuracy on retrospective test sets to causal validity under prospective deployment, from statistical fairness metrics to interventional equity guarantees, and from black-box pattern recognition to mechanistic interpretability verified against biological knowledge. Success requires coordinated advances across scalable causal discovery algorithms, validation frameworks suited to healthcare’s observational constraints, integration strategies for multi-modal multi-scale biological data, and regulatory pathways for systems making causal rather than purely associational claims. The path from proof-of-concept demonstrations to clinical implementation is neither straight nor simple, demanding collaboration between computational scientists, clinicians, biomedical researchers, and policymakers united by the recognition that healthcare AI’s next decade will be defined by which approaches achieve not just impressive metrics but genuine clinical trust through mechanistic understanding.
Causal graph neural networks address these fundamental limitations through principled causal inference frameworks (Fig. 1). By explicitly modelling causal structures within graph architectures, these methods identify invariant biological mechanisms rather than spurious correlations, enabling robust deployment across heterogeneous clinical settings and supporting mechanistic treatment optimisation through counterfactual reasoning.
@@ 2.1 Why biological systems demand graph-structured causal models
The mathematical foundation connecting causal inference to graph representations is based on Structural Causal Models (SCMs), which support a rigorous formalisation of cause-effect relationships [15]. An SCM $\mathcal{M}:=\langle U,V,F,P(U)\rangle$ specifies exogenous variables $U$ determined outside the model, endogenous variables $V$ defined within the model, structural equations $F=\{f_{1},\ldots,f_{m}\}$ where each $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and probability distribution $P(U)$ over exogenous variables [40]. This formulation naturally induces a directed acyclic graph $\mathcal{G}(\mathcal{M})$ where nodes represent variables and directed edges encode functional dependencies through the parent sets $\text{pa}(V_{i})$. The joint distribution factorises according to the Markov property: $P(V_{1},\ldots,V_{m})=\prod_{i=1}^{m}P(V_{i}|\text{pa}(V_{i}))$ [15]. Interventions in this framework correspond to “graph surgery,” modifying the causal graph by removing incoming edges to intervened variables and setting their values, fundamentally altering the generative process [15, 41]. This graph-theoretic representation of causation aligns precisely with the message-passing architecture of graph neural networks, enabling implementation of causal reasoning through differentiable neural operations.
@@ 2.2 From association to intervention to counterfactuals
Graph-structured causal models enable interventional prediction through Pearl’s do-calculus, which provides three rules for transforming interventional distributions into estimable observational quantities under specified graph conditions [15]. Rule 1 allows ignoring observations on variables that don’t affect the outcome through paths independent of the intervention. Rule 2 permits converting interventions to observations when appropriate conditional independence holds. Rule 3 enables removing interventions when they do not affect the outcome through certain graph paths. The interventional Variational Graph Auto-Encoder (iVGAE) implements these operations through differentiable neural layers, incorporating interventional transformations in both encoder and decoder networks to ensure that the learnt representations encode causal rather than associational relationships [41]. When the model identifies a connection between two proteins or two brain regions, it learns whether this connection would persist under interventions, whether it represents a causal pathway or a spurious correlation.
@@
The convergence of causal inference and graph learning thereby enables a transformative vision: patient-specific Causal Digital Twins, dynamic computational models where clinicians perform in silico experiments to test interventions before clinical administration. Such a system would integrate the multi-omics profile of a patient (genomics, transcriptomics, proteomics), longitudinal imaging that captures organ structure and function, clinical history including treatments and responses, and knowledge graphs encoding established biological mechanisms [30, 31]. The underlying causal GNN architecture would learn patient-specific parameterisations of general biological mechanisms, enabling simulation of therapeutic interventions through the do-operator framework by setting treatment variables, propagating effects through learnt causal pathways, and predicting outcomes at molecular, cellular, and phenotypic levels. For an oncology patient, the digital twin might simulate five drug combinations, predicting effects on specific oncogenic pathways, normal tissue toxicity, and progression-free survival, thereby identifying optimal personalised therapy before administering a single dose. For a neuropsychiatric patient, it could simulate neuromodulation protocols, predicting which brain circuits would be affected and how symptoms would respond. This vision, while ambitious, rests on methodological foundations now being established of GNN architectures for interventional prediction, counterfactual generation frameworks for exploring alternative scenarios, invariant learning methods ensuring robust generalisation, and multi-modal integration strategies combining diverse data modalities through shared causal structure [28, 46]. Figure 2 quantifies these theoretical distinctions, demonstrating how traditional graph neural networks can achieve comparable associational performance but exhibit substantial degradation for interventional and counterfactual queries essential for clinical decision-making.
[…]
@@
Box 1: Formal foundations of causal graph neural networks Structural Causal Models An SCM $\mathcal{M}:=\langle U,V,F,P(U)\rangle$ defines the data-generating process, where exogenous variables $U=\{U_{1},\ldots,U_{n}\}$ represent unmodelled background factors, endogenous variables $V=\{V_{1},\ldots,V_{m}\}$ are defined by structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and $P(U)$ governs exogenous randomness [15]. The induced causal graph $\mathcal{G}(\mathcal{M})$ contains directed edges $V_{j}\to V_{i}$ for each $V_{j}\in\text{pa}(V_{i})$. Interventions and the do-operator Intervention $\text{do}(X=x)$ modifies the SCM by replacing the structural equation for $X$ with $X:=x$, removing incoming edges in the graph. The post-intervention distribution $P(Y|\text{do}(X=x))$ differs from observational $P(Y|X=x)$ whenever confounding exists [15]. Graph neural network implementation GNNs implement message-passing: $\mathbf{h}_{i}^{(l+1)}=\sigma\left(\mathbf{W}^{(l)}\mathbf{h}_{i}^{(l)}+\sum_{j\in\mathcal{N}(i)}\mathbf{M}^{(l)}\mathbf{h}_{j}^{(l)}\right)$, where $\mathbf{h}_{i}^{(l)}$ is the representation of node $i$ at layer $l$, $\mathcal{N}(i)$ are neighbours, and $\mathbf{W},\mathbf{M}$ are learnable parameters. This architecture naturally aligns with causal graphs when edges represent causal relationships and message-passing implements causal influence propagation [41]. Causal identifiability $P(Y|\text{do}(X))$ is identifiable if it can be expressed in terms of the observational distribution $P(V)$ given the causal graph $\mathcal{G}$. The do-calculus provides three rules for such expressions [15]: $P(Y|\text{do}(X),Z,W)$ $=P(Y|\text{do}(X),W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}}}Z|X,W$ (1) $P(Y|\text{do}(X),\text{do}(Z),W)$ $=P(Y|\text{do}(X),Z,W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}\underline{Z}}}Z|X,W$ (2) $P(Y|\text{do}(X),\text{do}(Z),W)$ $=P(Y|\text{do}(X),W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}\overline{Z(W)}}}Z|X,W$ (3) where $\mathcal{G}_{\overline{X}}$ denotes the graph with edges into $X$ removed, $\mathcal{G}_{\underline{Z}}$ has edges out of $Z$ removed, and $\perp$ indicates d-separation. Counterfactual computation For unit $u$ and intervention $X=x$, the counterfactual outcome $Y_{x}(u)$ is computed via: (1) Abduction: infer $U=u$ from observed data; (2) Action: modify SCM to $\mathcal{M}_{x}$ with $X:=x$; (3) Prediction: evaluate $Y$ under $\mathcal{M}_{x}$ with $U=u$ [40]. This enables patient-specific predictions unavailable from population-level interventional distributions.

### v3

@@ Abstract
Healthcare artificial intelligence systems routinely fail when deployed across institutions, with documented performance drops and perpetuation of discriminatory patterns embedded in historical data. This brittleness stems, in part, from learning statistical associations rather than causal mechanisms. Causal graph neural networks address this triple crisis of distribution shift, discrimination, and inscrutability by combining graph-based representations of biomedical data with causal inference principles to learn invariant mechanisms rather than spurious correlations. This Perspective examines methodological foundations spanning structural causal models, disentangled causal representation learning, and techniques for interventional prediction and counterfactual reasoning on graphs. We analyse applications demonstrating clinical value across psychiatric diagnosis through brain network analysis, cancer subtyping via multi-omics causal integration, continuous physiological monitoring with mechanistic interpretation, and drug recommendation correcting prescription bias. These advances establish foundations for patient-specific Causal Digital Twins, enabling in silico clinical experimentation, with integration of large language models for hypothesis generation and causal graph neural networks for mechanistic validation. Substantial barriers remain, including computational requirements precluding real-time deployment, validation challenges demanding multi-modal evidence triangulation beyond cross-validation, and risks of causal-washing where methods employ causal terminology without rigorous evidentiary support. We propose tiered frameworks distinguishing causally-inspired architectures from causally-validated discoveries and identify critical research priorities, including hierarchical causal discovery exploiting biological modularity, integration of diverse data modalities, and regulatory pathways for systems making causal rather than purely associational claims. Success requires balancing theoretical ambition with empirical humility, computational sophistication with clinical interpretability, and transformative vision with uncompromising validation standards.
@@ 1 Introduction
Randomised controlled trials remain the gold standard for estimating causal effects, yet they address population-average treatment effects under specific protocols rather than patient-specific responses, cannot explore the combinatorial space of multi-drug interactions, and are infeasible for many clinically relevant questions due to ethical, temporal, or cost constraints. Causal graph methods extend this inferential capacity by enabling counterfactual reasoning for individual patients, leveraging observational data where randomisation is impossible, and scaling causal analysis to high-dimensional molecular and physiological networks beyond experimental reach.
@@
The convergence of causal inference with graph neural networks addresses this challenge. Recent theoretical advances in causal representation learning provide a path forward. By explicitly modelling causal structures within graph architectures, we can develop models that identify genuine therapeutic targets, predict treatment responses under interventions never observed in training data, and discover disease mechanisms rather than spurious correlations [28, 29]. This convergence is particularly timely given the four concurrent developments. First, the accumulation of multi-modal biomedical data, from single-cell sequencing that captures 20,000+ gene expression profiles to continuous physiological monitoring that generates terabytes of temporal dynamics, provides unprecedented resolution for causal discovery. Datasets now capture interventional outcomes and temporal dependencies that were previously unavailable [30, 31, 32]. Second, regulatory frameworks increasingly demand explainable AI in healthcare, which requires not merely predictions, but a mechanistic understanding that clinicians can verify against domain expertise and biological knowledge [33, 34]. Third, documented failures of correlation-based models under distribution shift have created urgency for methods achieving robust generalisation through causal invariance. Fourth, the integration of causal inference with fairness constraints enables rigorous identification and mitigation of discrimination transmitted along specific causal pathways, providing tools for distinguishing legitimate predictive factors from discriminatory proxies, a distinction central to anti-discrimination law yet absent from traditional machine learning approaches [13, 35].
@@
This Perspective examines causal graph neural networks, methods that integrate causal inference principles with graph-based deep learning architectures, as a methodological framework for transforming AI in healthcare from reactive pattern recognition to proactive causal medicine. Unlike previous reviews that addressed GNNs in healthcare [36, 37] or causal inference in medicine [16, 38] separately, we focus specifically on their integration, methods in which both graph structure and causal reasoning prove essential for clinical impact. We analyse how causal principles enhance GNN architectures for healthcare tasks, moving beyond correlation to mechanism, and examine current applications that demonstrate clinical utility across psychiatric diagnosis, cancer treatment, physiological monitoring, and drug discovery. We also identify fundamental challenges in scaling these methods to real-world deployment, including computational complexity, validation without ground truth, fairness under distribution shift, and regulatory considerations. Our analysis reveals both transformative promise and substantial barriers: while CIGNNs achieve superior generalisation across healthcare institutions, identify biologically plausible mechanisms, and maintain fairness properties under distribution shift, they face computational requirements precluding real-time application, difficulty validating causal claims from observational data, and gaps between methodological sophistication and clinical interpretability.
@@
The result of this research is the development of patient-specific Causal Digital Twins, dynamic computational models built on causal GNN frameworks that enable clinicians to perform in silico experiments before clinical intervention. The digital twin paradigm has emerged as a transformative framework for personalised medicine, offering dynamic computational representations of individual patients that evolve through continuous data integration [39, 40]. Recent conceptualisation identifies four essential capabilities that distinguish medical digital twins from conventional simulation models: explainability through mechanistic rather than black-box architectures, intervenability enabling virtual testing of therapeutic scenarios, learnability allowing patient-specific model refinement over time, and diversability supporting uncertainty quantification through ensemble approaches [39]. Early clinical successes, notably the artificial pancreas for type 1 diabetes management, demonstrate feasibility within circumscribed physiological systems [41, 42]. However, extending digital twins to complex multisystem diseases requires methodological foundations that current data-driven approaches cannot provide. Causal graph neural networks could supply the missing theoretical infrastructure. Structural causal models enable the interventional reasoning essential for virtual treatment testing, counterfactual computation supports patient-specific outcome prediction, and invariant causal learning can potentially ensure robust generalisation across healthcare contexts. The integration of causal inference with graph-based patient modelling thus can allow us to transform digital twins from conceptual aspiration to methodologically grounded clinical tools. Imagine a clinician treating advanced cancer who could load the patient’s multi-omics profile, brain imaging, and clinical history into such a system, then simulate five drug combinations to predict effects on specific tumour pathways, toxicity risks, and progression-free survival, identifying optimal personalised therapy before administering a single dose. This vision, while ambitious, rests on methodological foundations now being established through causality-inspired architectures for interventional prediction, counterfactual generation, and robust multi-modal integration. Realising this potential requires not only technical innovation but fundamental rethinking of how we evaluate and deploy AI in medicine. We need to shift from predictive accuracy on retrospective test sets to causal validity under prospective deployment, from statistical fairness metrics to interventional equity guarantees, and from black-box pattern recognition to mechanistic interpretability verified against biological knowledge. Success requires coordinated advances across scalable causal discovery algorithms, validation frameworks suited to healthcare’s observational constraints, integration strategies for multi-modal multi-scale biological data, and regulatory pathways for systems making causal rather than purely associational claims. The path from proof-of-concept demonstrations to clinical implementation is neither straight nor simple, demanding collaboration between computational scientists, clinicians, biomedical researchers, and policymakers united by the recognition that healthcare AI’s next decade will be defined by which approaches achieve not just impressive metrics but genuine clinical trust through mechanistic understanding.
Causal graph neural networks address these fundamental limitations through principled causal inference frameworks (Fig. 1). By explicitly modelling causal structures within graph architectures, these methods identify invariant biological mechanisms rather than spurious correlations, enabling robust deployment across heterogeneous clinical settings and supporting mechanistic treatment optimisation through counterfactual reasoning. This Perspective proceeds in three stages. Section 2 establishes the theoretical foundations connecting structural causal models to graph neural network architectures. Section 3 presents methodological advances organised by clinical capability, disentanglement, intervention prediction, counterfactual generation, robustness, and fairness, providing the technical building blocks. Section 4 demonstrates clinical impact, showing how these methods translate to diagnosis, prognosis, treatment selection, and monitoring across healthcare domains. Section 5 synthesises these advances toward Causal Digital Twins, articulating how the methodological components combine to enable patient-specific in silico experimentation. Section 6 addresses the substantial barriers, computational, validation, and translational, that must be overcome for clinical deployment. We note that Sections 2–4 primarily synthesise existing methods and empirical evidence, while Section 5 articulates an aspirational framework, Causal Digital Twins, whose full realisation remains prospective. Section 6 addresses the substantial methodological, computational, and translational barriers separating current capabilities from this vision.
@@ 2.1 Why biological systems demand graph-structured causal models
The mathematical foundation connecting causal inference to graph representations is based on Structural Causal Models (SCMs), which support a rigorous formalisation of cause-effect relationships [15]. An SCM $\mathcal{M}:=\langle U,V,F,P(U)\rangle$ specifies exogenous variables $U$ determined outside the model, endogenous variables $V$ defined within the model, structural equations $F=\{f_{1},\ldots,f_{m}\}$ where each $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$ with $\text{pa}(V_{i})$ denoting the parent set of $V_{i}$, comprising all variables that directly cause $V_{i}$ in the causal graph, and probability distribution $P(U)$ over exogenous variables [44]. This formulation naturally induces a directed acyclic graph $\mathcal{G}(\mathcal{M})$ where nodes represent variables and directed edges encode functional dependencies through the parent sets $\text{pa}(V_{i})$. The joint distribution factorises according to the Markov property: $P(V_{1},\ldots,V_{m})=\prod_{i=1}^{m}P(V_{i}|\text{pa}(V_{i}))$ [15]. Interventions in this framework correspond to “graph surgery,” modifying the causal graph by removing incoming edges to intervened variables and setting their values, fundamentally altering the generative process [15, 45]. This graph-theoretic representation of causation aligns precisely with the message-passing architecture of graph neural networks, enabling implementation of causal reasoning through differentiable neural operations.
A critical modelling choice underlying any causal graph concerns the semantic definition and resolution of nodes [46, 47]. In biomedical applications, nodes may represent entities at vastly different levels of abstraction such as individual genes versus pathways, single neurons versus brain regions, specific metabolites versus metabolic modules. This granularity is not merely a technical convenience but fundamentally determines which causal relationships can be identified and what interventions the graph implies [48]. Causal structure discovered at one resolution may not transfer to finer or coarser scales, as relationships that hold between brain regions might decompose into opposing sub-regional effects, while molecular interactions may aggregate into emergent pathway-level causation invisible at the gene level. Throughout this Perspective, we explicitly note the resolution at which different methods operate, as this choice carries direct consequences for both the biological interpretation of discovered causal structures and the experimental designs capable of validating them.
A related question concerns the epistemic status of existing biological network representations. Gene regulatory networks encode regulatory relationships that carry causal interpretation, perturbing transcription factor X should alter expression of target gene Y. However, GRNs inferred from observational expression data may conflate correlation with causation, with edges reflecting shared upstream regulators or cellular state confounding rather than direct regulation. Recent large-scale benchmarking on single-cell perturbation data reveals that even state-of-the-art inference methods struggle to recover true causal relationships when validated against interventional ground truth [49]. Systems biology models, comprising ordinary differential equations encoding metabolic and signalling dynamics, occupy different epistemic ground. These models encode mechanistic hypotheses derived from biochemical knowledge, enzyme kinetics, binding affinities, reaction stoichiometry, and thus represent causal mechanisms more explicitly. Yet parameters fitted to observational data may encode environment-specific rather than invariant relationships. Protein-protein interaction networks used by methods such as PDGrapher as ”causal graph approximations” [50] contain both direct causal interactions and indirect associations through protein complexes. We advocate treating biological networks as candidate causal structures requiring validation through interventional evidence, not established ground truth. The CiGNN framework exemplifies this by using causal discovery algorithms to identify genuinely causal physiological relationships rather than assuming database networks encode causation [51].
@@ 2.2 From association to intervention to counterfactuals
Graph-structured causal models enable interventional prediction through Pearl’s do-calculus, which provides three rules for transforming interventional distributions into estimable observational quantities under specified graph conditions (Box 1) [15]. Rule 1 allows ignoring observations on variables that don’t affect the outcome through paths independent of the intervention. Rule 2 permits converting interventions to observations when appropriate conditional independence holds. Rule 3 enables removing interventions when they do not affect the outcome through certain graph paths. The interventional Variational Graph Auto-Encoder (iVGAE) implements these operations through differentiable neural layers, incorporating interventional transformations in both encoder and decoder networks to ensure that the learnt representations encode causal rather than associational relationships [45]. When the model identifies a connection between two proteins or two brain regions, it learns whether this connection would persist under interventions, whether it represents a causal pathway or a spurious correlation.
@@
Throughout this Perspective, we distinguish methods by the strength of their causal claims: causally-inspired architectures that employ causal concepts without explicit causal claims, causally-structured methods that estimate causal quantities under stated assumptions, and causally-validated approaches whose claims are corroborated through multi-modal evidence (see Section 6.3 for formal definitions).
@@
Patient-specific Causal Digital Twins, dynamic computational models where clinicians perform in silico experiments to test interventions before clinical administration could integrate the multi-omics profile of a patient (genomics, transcriptomics, proteomics), longitudinal imaging that captures organ structure and function, clinical history including treatments and responses, and knowledge graphs encoding established biological mechanisms [30, 31]. The underlying causal GNN architecture would learn patient-specific parameterisations of general biological mechanisms, enabling simulation of therapeutic interventions through the do-operator framework by setting treatment variables, propagating effects through learnt causal pathways, and predicting outcomes at molecular, cellular, and phenotypic levels. For an oncology patient, the digital twin might simulate five drug combinations, predicting effects on specific oncogenic pathways, normal tissue toxicity, and progression-free survival, thereby identifying optimal personalised therapy before administering a single dose. For a neuropsychiatric patient, it could simulate neuromodulation protocols, predicting which brain circuits would be affected and how symptoms would respond. This vision, while ambitious, rests on methodological foundations now being established of GNN architectures for interventional prediction, counterfactual generation frameworks for exploring alternative scenarios, invariant learning methods ensuring robust generalisation, and multi-modal integration strategies combining diverse data modalities through shared causal structure [28, 55]. Figure 2A illustrates these theoretical distinctions, depicting the conceptual framework through which traditional graph neural networks may achieve comparable associational performance but are expected to exhibit substantial degradation for interventional and counterfactual queries essential for clinical decision-making.
@@
Box 1: Formal foundations of causal graph neural networks Structural Causal Models An SCM $\mathcal{M}:=\langle U,V,F,P(U)\rangle$ defines the data-generating process, where exogenous variables $U=\{U_{1},\ldots,U_{n}\}$ represent unmodelled background factors, endogenous variables $V=\{V_{1},\ldots,V_{m}\}$ are defined by structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and $P(U)$ governs exogenous randomness [15]. The induced causal graph $\mathcal{G}(\mathcal{M})$ contains directed edges $V_{j}\to V_{i}$ for each $V_{j}\in\text{pa}(V_{i})$. Interventions and the do-operator Intervention $\text{do}(X=x)$ modifies the SCM by replacing the structural equation for $X$ with $X:=x$, removing incoming edges in the graph. The post-intervention distribution $P(Y|\text{do}(X=x))$ differs from observational $P(Y|X=x)$ whenever confounding exists [15]. Graph neural network implementation GNNs implement message-passing: $\mathbf{h}_{i}^{(l+1)}=\sigma\left(\mathbf{W}^{(l)}\mathbf{h}_{i}^{(l)}+\sum_{j\in\mathcal{N}(i)}\mathbf{M}^{(l)}\mathbf{h}_{j}^{(l)}\right)$, where $\mathbf{h}_{i}^{(l)}$ is the representation of node $i$ at layer $l$, $\mathcal{N}(i)$ are neighbours, and $\mathbf{W},\mathbf{M}$ are learnable parameters. This architecture naturally aligns with causal graphs when edges represent causal relationships and message-passing implements causal influence propagation [45]. Causal identifiability $P(Y|\text{do}(X))$ is identifiable if it can be expressed in terms of the observational distribution $P(V)$ given the causal graph $\mathcal{G}$. The do-calculus provides three rules for such expressions [15]: $P(Y|\text{do}(X),Z,W)$ $=P(Y|\text{do}(X),W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}}}Z|X,W$ (1) $P(Y|\text{do}(X),\text{do}(Z),W)$ $=P(Y|\text{do}(X),Z,W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}\underline{Z}}}Z|X,W$ (2) $P(Y|\text{do}(X),\text{do}(Z),W)$ $=P(Y|\text{do}(X),W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}\overline{Z(W)}}}Z|X,W$ (3) where $\mathcal{G}_{\overline{X}}$ denotes the graph with edges into $X$ removed, $\mathcal{G}_{\underline{Z}}$ has edges out of $Z$ removed, and $\perp$ indicates d-separation. Rule 1 permits ignoring observations on variables that do not affect the outcome through paths independent of the intervention; Rule 2 allows converting interventions to observations when appropriate conditional independence holds; Rule 3 enables removing interventions when they do not influence the outcome through certain graph paths. Counterfactual computation For unit $u$ and intervention $X=x$, the counterfactual outcome $Y_{x}(u)$ is computed via: (1) Abduction: infer $U=u$ from observed data; (2) Action: modify SCM to $\mathcal{M}_{x}$ with $X:=x$; (3) Prediction: evaluate $Y$ under $\mathcal{M}_{x}$ with $U=u$ [44]. This enables patient-specific predictions unavailable from population-level interventional distributions.
Figure 2: Conceptual framework illustrating theoretical advantages of causal over associational inference across Pearl’s causal hierarchy. A, Theoretical dissociation between associational and interventional performance predicted by causal hierarchy theory. The “causal gap” quantifies expected performance loss when models trained on observational data are applied to interventional queries, reflecting the fundamental limitation that associational patterns may not persist under intervention. B, Idealised robustness to simulated unmeasured confounding. Conceptual curves illustrate the theoretical expectation that traditional methods degrade more steeply as unmeasured confounder strength increases, while causal architectures that explicitly model confounding structure are expected to maintain robustness. Confounding strength parameterised as maximum risk ratio for unmeasured confounder associations with treatment and outcome following sensitivity analysis frameworks [56]. Shaded regions represent idealised uncertainty bands illustrating expected variance patterns. Causally-Inspired GNN (Tier 1) denotes architectures employing causal concepts such as invariance regularisation to improve robustness without explicit causal claims. Causally-Validated GNN (Tier 3) denotes methods whose causal claims are corroborated through multi-modal evidence triangulation. See Section 6.3 for formal tier definitions.